OpenStudy (anonymous):

(2x-2y)(x+2y) Help Please, anyone !

OpenStudy (psymon):

Would you know the order in which you foil this? Or are we not sure about that either?

OpenStudy (anonymous):

Nope, I Tried But Its Harder That Way To Me.

OpenStudy (psymon):

Well, let me know if what I do is still confusing. So basically, the 2x in the left parenthesis has to multiply by both things in the right. And also, the -2y has to multiply both things in the right parenthesis. So I'll stack my operations so they look a bit better visually. So the first operation I do is 2x * x = 2x^2 next 2x * 2y = 4xy next -2y * x = -2xy lastly -2y*2y = -4y^2 Now I add up all these answers: 2x^2 + 4xy -2xy - 4y^2. The two middle terms are like terms, so I combine them and end up with a final answer of 2x^2 + 2xy - 4y^2

OpenStudy (anonymous):

THanks I Have Another One (s+2t^2)(2s-2t^2)

OpenStudy (psymon):

You think you would be able to try and multiply things out in the same order and see if you get the right answer?

OpenStudy (anonymous):

I Know How To Do It But I End Up GEtting Confused And Then I Get The Wrong Answer.

OpenStudy (psymon):

Can you maybe try and type it out the same way I did? Then I can point out what you may be doing wrong. You'll have to eventually do it on your own anyway, so let's see what youre doing wrong and maybe we can fix the error.

OpenStudy (okra77):